Sally is 10 years older than Tim. In 8 years, the sum of their ages will be 40. Determine how old Sally and Tim are now.
Answer by lynnlo(4176)

SALLY,,,,,,,,,,,,,,TIM
SALLY==A
TIM====B
A=B+10
A+8+B+7=40
A+B=40-10
A+B=30
A-B=10
A+B=30
A-B=10
2A=40
A=20
SALLY==========40
TIM============20